bubbly1belle:

When trying to solve y=2/4x+6 and y is 5, what is x?

jhonyy9:

like a first step \[y = \frac{ 2 }{ 4 }x +6\] \[y = \frac{ 1 }{ 2 }x +6\] given y=5 and now just substitute the 5 in place of y inside equation \[5 = \frac{ 1 }{ 2 }x +6\] multiplie both sides by 2 and will get \[10 = x +12\] \[x = 10-12\] \[x=-2\]